OpenStudy (anonymous):

The perimeter of a rectangle is 30ft. the difference between the length and width is 1ft. find the width

OpenStudy (anonymous):

perimeter P=2L+2W=30, L-W=1ft, L=1+W 2L+2W=30 2(1+W)+2W=30 2+2W+2W=30 4W=28 W=28/4=7 ans